OpenStudy (anonymous):

You drop a ball from a height of 0.5 meter. Each curved path has 52% of the height of the previous path. a. Write a rule for the sequence using centimeters. The initial height is given by the term n = 1. b. What height will the ball be at the top of the third path? a. A(n) = 0.5 * (0.52)^n-1; 0.14 cm b.A(n) = 0.52 * (0.5)^n-1 ; 0.13 cm c.A(n) = 50 * (52)^n-1 ; 135,200 cm d.A(n) = 50 * (0.52)^n-1 ; 13.52 cm

OpenStudy (anonymous):

its d

OpenStudy (nurali):

A(n)=50(52/100)^n-1 Convert (0.5m=50cm) n=3 A(3)=50(0.52)^3-1 A(3)=50(0.52)^2 A(3)=50(0.2704) A(3)=13.52 cm

OpenStudy (anonymous):

thank you guys

OpenStudy (nurali):

Anytime.